OpenStudy (anonymous):

Simplify (12q^3r^2)(4qr^4)

OpenStudy (anonymous):

You need to know the index rule: \[a^m\times a^n=a ^{m+n}\] In your case: \[12q^3\times r^2\times 4q\times r^4\] Group all the Qs together and all the Rs together and apply the index rule

OpenStudy (anonymous):

so the answer would be 48q^4r^6?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

okay, thanks.

OpenStudy (anonymous):

@SamSmiths Something a bit ambiguous in the question: is it 4qxr^4 or (4qr)^4?

OpenStudy (anonymous):

4q*r^4

OpenStudy (anonymous):

Yes then your answer is correct :)